OpenStudy (anonymous):

Solve the following system using the substitution method. 16x + 3y = 0 8x + y = 1 Solve the following system using the substitution method. 16x + 3y = 0 8x + y = 1 @Mathematics

OpenStudy (eyust707):

first solve for y.. basically take one of the equations and make it look like y= ???? then use that ??? as y in the other equation

OpenStudy (anonymous):

i don't know what you mean by that.

OpenStudy (anonymous):

solve for y in the second equation..it is : y=-8x+1 and then just take that substitute back into the first equation for 'y' and solve for 'x'

OpenStudy (anonymous):

if i give y'all the answers, can you just tell me the right one?

OpenStudy (eyust707):

lol but then u dont learn anything =/

OpenStudy (anonymous):

SURE I can help with that.

OpenStudy (anonymous):

yeaa eyust is right about that as well.

OpenStudy (anonymous):

i wouldn't learn anything with you guys telling me how to do it, i suck at math. (:

OpenStudy (anonymous):

16x + 3y = 0 8x + y = 1 Take the second equation and isolate the y y = -8x + 1 The substitute the y in the first equation with -8x + 1 16x + 3(-8x + 1) = 0 16x - 24x + 3 = 0 -8x = -3 /-8 /-8 x = 3/8 The plug in the x for the second equation 8(3/8) + y = 1 24/8 + y = 1 3 + y = 1 -3 -3 y = -2 Then check your work, I hope those are the right answers

OpenStudy (anonymous):

its not that hard...as mito showed it to you.

OpenStudy (eyust707):

savanna do you have skype?

OpenStudy (anonymous):

yeah..?

OpenStudy (eyust707):

skype me ill help ya out..

OpenStudy (eyust707):

my names the same as on here

OpenStudy (anonymous):

mkay, i'll do it in a few.

OpenStudy (eyust707):

kk ill prolly be on for the next hour or so